# A Note on Optimal Product Pricing

### v2

@@ 2.1 Prices and profit
the (logarithmic) fractional demand change, with respect to the nominal demand. We let $\delta\in{\mbox{\bf R}}^{n}$ denote the vector of demand changes.
@@
where $E\in{\mbox{\bf R}}^{n\times n}$ is the elasticity matrix, also referred to as the Slutsky substitution matrix. The entry $E_{ij}$ is the elasticity of the demand for product $i$ with respect to the price of product $j$. When $i=j$, this is called a self-elasticity. When $i\neq j$, this is called a cross-elasticity.
We make no assumptions about the elasticity matrix, but we mention here some typical attributes; see, e.g., [Var92, MCWG95] for more on elasticity matrices. In almost all practical cases, the self-elasticities $E_{ii}$ are negative, which means that an increase in price results in a decrease in demand for that product. Cross-elasticities can be positive or negative. For example, when products $i$ and $j$ are substitutes for each other, $E_{ij}$ and $E_{ji}$ will be positive, as a higher price of one will result in higher demand for the other product (which will be bought as a substitute). When two products are complements (e.g., printer and ink), then their cross-elasticity will be negative. Using the example of printer and ink, increasing the price for ink will decrease the demand for ink, and the demand for printing in general. While the order of magnitude of the self-elasticities is typically 1-1, the cross-elasticities are typically smaller in magnitude. The elasticity matrix is typically sparse, e.g., block diagonal, with the blocks representing similar or related products. In general, it need not be symmetric [KSCdPM00, MCWG95, §2.F].
Profit.

### v3

@@ 1.1 Optimal pricing
There is related work on the theory of multi-product pricing [AV18], which focuses on economic models, but not on numerical optimization methods for choosing prices subject to constraints, as we do. For example, it is common to model demand as the maximizer of the expenditure-adjusted or expenditure-constrained utility of purchasing a portfolio of products, which is sometimes referred to as Marshallian demand [AV18, MCWG95, §3.D]. A closely related concept is Hicksian demand, which minimizes the total expenditure when purchasing a portfolio of products, subject to a minimum utility [LP09]. Further, multi-product pricing is studied in the context of monopolies [Oi71, BM82, AJPT16], and Cournot oligopolies, where a few firms control a market and maximize their respective profits in a game-theoretic sense [Viv99, JM06, NS18]. In this paper, however, we take a simple model of demand, based on an elasticity matrix or Slutsky substitution matrix, as described in, e.g., [MCWG95, §2.F].
@@ 2.2 Demand
the (logarithmic) fractional demand change, with respect to the nominal demand. We let $\delta\in{\mbox{\bf R}}^{n}$ denote the vector of demand changes. (It follows that the revenue for product $i$ is $p_{i}d_{i}$, and the nominal revenue for product $i$ is $p_{i}^{\text{nom}}d_{i}^{\text{nom}}$, denoted by $r_{i}^{\text{nom}}$.)
@@
where $E\in{\mbox{\bf R}}^{n\times n}$ is the elasticity matrix, also referred to as the Slutsky substitution matrix. The entry $E_{ij}$ is the elasticity of the demand for product $i$ with respect to the price of product $j$. When $i=j$, this is called a self-elasticity. When $i\neq j$, this is called a cross-elasticity. This model is basically the linearization of $\delta$ as a function of $\pi$, around $\pi=0$ (i.e., nominal prices) [Per09, Var14].
We make no assumptions about the elasticity matrix, but we mention here some typical attributes; see, e.g., [Var92, MCWG95] for more on elasticity matrices. In almost all practical cases, the self-elasticities $E_{ii}$ are negative, which means that an increase in price results in a decrease in demand for that product. Cross-elasticities can be positive or negative. For example, when products $i$ and $j$ are substitutes for each other, $E_{ij}$ and $E_{ji}$ will be positive, as a higher price of one will result in higher demand for the other product (which will be bought as a substitute). When two products are complements (e.g., printer and ink), then their cross-elasticity will be negative. Using the example of printer and ink, increasing the price for ink will decrease the demand for ink, as well as the demand for printers. While the order of magnitude of the self-elasticities is typically 1-1, the cross-elasticities are typically smaller in magnitude. The elasticity matrix is typically sparse, e.g., block diagonal, with the blocks representing similar or related products.
Utility-based demand.
We mention here a different but related demand model, based on a utility function. A utility-based demand model has the form
[…]
where $U:{\mbox{\bf R}}_{+}^{n}\rightarrow{\mbox{\bf R}}$ is a strictly concave and increasing utility function; see, e.g., [AV18]. This means that the demand maximizes the utility of purchasing a portfolio of products, minus the total expenditure.
The utility-based demand (2) does not in general have the form (1), but we can form a local approximation that does. We first form the first order Taylor approximation of $\mathcal{D}$ at the nominal price and demand, to obtain
$d\approx d^{\text{nom}}+D\mathcal{D}(p^{\text{nom}})(p-p^{\text{nom}}),$
where $D\mathcal{D}$ is the derivative of $\mathcal{D}$, which can be shown to have the form
$D\mathcal{D}(p^{\text{nom}})=\left(\nabla^{2}U(p^{\text{nom}})\right)^{-1},$
assuming $U$ is twice differentiable. From this we obtain the first order elasticity approximation $\delta\approx E^{\text{util}}\pi$, with
$E^{\text{util}}=\mathop{\bf diag}(d^{\text{nom}})^{-1}\left(\nabla^{2}U(p^{\text{nom}})\right)^{-1}\mathop{\bf diag}(p^{\text{nom}}),$
where we use the first order approximations $\delta_{i}\approx(d_{i}-d_{i}^{\text{nom}})/d_{i}^{\text{nom}}$ and $\pi_{i}\approx(p_{i}-p_{i}^{\text{nom}})/p_{i}^{\text{nom}}$. Since $\left(\nabla^{2}U(p^{\text{nom}})\right)^{-1}$ is symmetric and negative definite, it follows that if the elasticity matrix $E$ comes from linearization of a utility-based demand model, then $\mathop{\bf diag}(r^{\text{nom}})E$, i.e., $E$ with its rows scaled by the nominal revenues, is symmetric and negative definite. This implies, for example, that $E_{ii}<0$.
Here we are simply observing that if the elasticity matrix comes from a utility-based demand, then it has this specific form; in the sequel, however, we make no assumptions about the structure of $E$.
2.3 Profit
@@ 5.2 Results
Acknowledgments
The authors thank Garrett van Ryzin for his insightful discussions and valuable suggestions.